# Exploring Holomorphic Retracts

### v6

@@ Conclusion
which is negative as $q<1$. We conclude therefore that $D_{q}$ is neither convex nor concave (though of-course pseudoconvex) at $(x,y)\in\partial D_{q}$. Owing to already made remarks this holds at each point of $\partial D_{q}$ other than the negligible set of non-smooth boundary points of $D_{q}$.
@@
Proceeding now towards the retracts of $D_{q}$ (where we no longer assume $N=2$ for that would be an oversimplification unlike the matter in the preceding paragraph), it must first be noted that by proposition 1.4, every holomorphic retract of $\ell_{q}$-ball through the origin is actually linear even when $q<1$; for $1$-dimensional retracts, this follows from theorem 1.3 but as highlighted in the introduction, the thrust here is on retracts of arbitrary dimension. So, if $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Of-course, we are only interested in non-trivial retracts, so $Y_{q}$ is of the form $Y_{q}=D_{q}\cap Y$ for some $r$-dimensional $\mathbb{C}$-linear subspace $Y$ of $\mathbb{C}^{N}$. We only need to show that $Y$ cannot be of the form
@@
$t_{j}=\frac{1}{\left(1+\sum_{k=r+1}^{N}|a_{j,k}|^{q}\right)^{1/q}}.$
@@
Since $P$ maps $\mathbb{C}^{N}$ into $Y$, $P(v_{j})\in Y$. Write
@@
Note that $P(v_{j})\in\overline{D_{q}}$ for all $j$ So, we get the following inequality for $1\leq j\leq r$,
@@
By reverse triangle inequality for $q^{\text{th}}$ power where $10<q<$ 1: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.4) becomes
$-\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~{}~{}k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
@@
$-\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~{}k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,j}\right|^{q}\right)+\sum_{1\leq m\leq r~{}~{}m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}\left|a_{j,l}~{}a_{k,s}u_{k,l}\right|^{q}\leq 0.$
@@
To make some noteworthy remarks here, let us also record a special case about all non-trivial retracts through the origin of such a domain in $\mathbb{C}^{2}$ where this result is not new (though our proof is direct and more elementary).
Proposition 5.2.
Let $q$ be an arbitrary positive number less than one. Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<1\}$ and the subspace $Y=\{(z,w)\in\mathbb{C}^{2}:z-mw=0\}$ (where $0\neq m\in\mathbb{C}$) of $\mathbb{C}^{2}$. Then the one dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$. In-fact, the holomorphic retracts through the origin are linear and more precisely, the intercepts of the coordinate axes with $D_{q}$.
Owing to remarks made in the introduction (in greater detail, the work of section 8 below), it follows as an immediate corollary that $D_{q}$ is not biholomorphic to any product domain; it must also be noted here that retracts can be used to immediately show inequivalence with convex domains – for instance, the non-existence of retracts of $D_{q}$ in certain directions suffices to prove its biholomorphic distinctness from bounded convex domains (the unbounded case is ruled out simply by hyperbolicity of $D_{q}$). A more important relevance of this example is noted next.
@@
We recall from [Vigue] that fixed point sets of holomorphic self-maps of a any given bounded (not necessarily balanced) convex domain $D$ are precisely the same as its retracts. The example of the above proposition proves that this coincidence fails as soon as we drop convexity. Indeed, for the non-convex (but pseudoconvex) balanced domain $D_{q}$ in $\mathbb{C}^{2}$ as above, the fixed point set of the holomorphic self-map of $D_{q}$ as simple as the flip $(z,w)\to(w,z)$, namely the linear subspace $\{(z,w)\in\mathbb{C}^{2}\;:\;w=z\}$, is not realizable as a holomorphic retract of $D_{q}$, which is part of the content of the proposition 5.2 and its higher dimensional analogue in theorem 1.9 above.
The $\ell_{q}$-balls dealt with above are bounded by holomorphically extreme real hypersurfaces. We next discuss examples for the ‘other extreme’ as well as some intermediate cases. We begin with an example where we can show existence of higher dimensional retracts; the purpose being: to record atleast one simple example to affirmatively answer the question as to whether there exists non-homogeneous pseudoconvex domains in $\mathbb{C}^{N}$ (i.e., when $N$ is not restricted to be $2$ or the domain being convex) with the property that there are retracts through every point of the domain and all possible dimensions; given Lempert’s apt remarks about the non-existence of retracts of dimension any bigger than one, in general. As our purpose is in recording the simplest cases first, the examples may well not be impressive.

### v7

@@ Conclusion
which is negative as $q<1$. We conclude therefore that $D_{q}$ is neither convex nor concave (though of course pseudoconvex) at $(x,y)\in\partial D_{q}$. Owing to already made remarks this holds at each point of $\partial D_{q}$ other than the negligible set of non-smooth boundary points of $D_{q}$.
Proposition 5.7.
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $L$ is a linear subspace of $\mathbb{C}^{N}$ such that ${\rm co}(\overline{D_{L}})\neq C_{L}$, where $C_{L}:=C\cap L$ with $C={\rm co}(\overline{D})$, denoting the convex hull of $\overline{D}$. Then $D_{L}:=D\cap L$ is not a linear retract of $D$.
Proof.
To prove this proposition by contradiction, assume that $D_{L}$ is a linear retract of $D$. Then there exists a linear projection $P:\mathbb{C}^{N}\to L$ such that $P$ maps $D$ onto $D_{L}$. Since ${\rm co}(\overline{D_{L}})\neq C_{L}$ by hypothesis, there exists a point $p$ lying within the convex set $C_{L}$ but not in ${\rm co}(\overline{D_{L}})$. Write $p\in C_{L}\subset C$ as the convex combination of extreme points of $C$ i.e. $p=\sum_{j=1}^{k}\alpha_{j}e_{j}$, (where $\alpha_{j}$’s are positive reals with $\sum_{j=1}^{k}\alpha_{j}=1$), where $e_{j}$’s are extreme points of $C$ (here of course $e_{j}$’s need not be the standard basis vectors, though we use this notation only in this proof for convenience and the extreme points of $C$ need not be finite in number). Such an expression is possible due to the Krein-Milman theorem applied to the compact convex set $C$. By the linearity of $P$, we have
$p=P(p)=\sum_{j=1}^{k}\alpha_{j}P(e_{j}).$
As $P(e_{j})\in\overline{D_{L}}$, this shows that $p$ can be written as the convex combination of points in $\overline{D_{L}}$. This implies that $p\in{\rm co}(\overline{D_{L}})$, which contradicts the hypothesis that ${\rm co}(\overline{D_{L}})\neq C_{L}$. ∎
We now proceed to prove theorem 1.10.
Proof of theorem 1.10:.
Assume that $L$ is not spanned by vectors only from $\partial D\cap\partial C$. We want to show that $D_{L}$ is not a linear retract of $D$. By the above proposition, it suffices to show that ${\rm co}(\overline{D_{L}})\neq C_{L}$. Now we claim that there exists a vector $v\notin\partial D\cap\partial C$ such that $v$ is an extreme point of ${\rm co}(\overline{D_{L}})$. To prove this claim by contradiction, assume that every extreme point of ${\rm co}(\overline{D_{L}})$ lies in $\partial D\cap\partial C$. We may write each $w\in{\rm co}(\overline{D_{L}})$, as a convex combination of the extreme points in ${\rm co}(\overline{D_{L}})$. Let $w^{\prime}$ be an arbitrary point in $L$. Note that there exists $\lambda\in\mathbb{C}^{*}$ such that $\lambda w^{\prime}\in D_{L}$. So, we may express $\lambda w^{\prime}$ as a convex combination of the extreme points of ${\rm co}(\overline{D_{L}})$. Since these extreme points all lie within $\partial D\cap\partial C$, we get that $w^{\prime}$ belongs to span of vectors in $\partial D\cap\partial C$. Since this holds for every $w^{\prime}\in L$, $L$ is spanned by vectors in $\partial D\cap\partial C$, which contradicts the hypotheses of the theorem. This finishes the proof of the claim.
By theorem 2.10.15 in [Meggn], $v\in\partial D_{L}=L\cap\partial D$. As $v\notin\partial C\cap\partial D$, we have $v\notin\partial C_{L}$. To prove ${\rm co}(\overline{D_{L}})\neq C_{L}$, choose $t_{0}>1$ such that $t_{0}v\in\partial C_{L}$. We want to show that $t_{0}v\notin{\rm co}(\overline{D_{L}})$. To prove this by contradiction, assume that $t_{0}v\in{\rm co}(\overline{D_{L}})$. Write $t_{0}v=c_{1}v_{1}+\ldots+c_{k}v_{k}$ (where $c_{j}$’s are positive reals with $\sum_{j=1}^{k}c_{j}=1$), where $v_{j}$’s are extreme points of ${\rm co}(\overline{D_{L}})$. After dividing $t_{0}$ on both sides, we obtain that
$v=c_{1}\left(\frac{v_{1}}{t_{0}}\right)+\ldots+c_{k}\left(\frac{v_{k}}{t_{0}}\right)$
Since $v_{j}/t_{0}\in D_{L}$ for each $j$, $v$ can be expressed as a convex combination of points in $D_{L}$. This contradicts the fact that $v$ is an extreme point of ${\rm co}(\overline{D_{L}})$. This proves that ${\rm co}(\overline{D_{L}})\neq C_{L}$. ∎
Lemma 5.8.
Let $D$ be a bounded pseudoconvex domain in $\mathbb{C}^{N}$ and $p$ a boundary point of $D$ near which $\partial D$ is smooth. Fix $p\in\partial D$ and suppose that the (real) affine line $l(p,v):=\{p+tv:t\in\mathbb{R}\}$ with $v\in T_{p}^{\mathbb{C}}(\partial D)$ does not intersect $D$. Then the real Hessian $\mathcal{H}r(p)$ along $v$ is non-negative, where $r$ is any smooth defining function near $p$.
Proof.
Let us begin with the standard parametrization of the (real) line $l$ passing through $p$ in the direction of $v$ with $v\in T_{p}^{\mathbb{C}}(\partial D)$, namely $\varphi(t)=p+tv$. Let $r$ be a local defining function of $D$ defined on an open neighbourhood $U$ of $p$. So, $D\cap U=\{z\in U:r(z)<0\}$. Define $V:=\varphi^{-1}(U)$ and consider the smooth function $\psi:V\to\mathbb{R}$ defined by $\psi(t)=r\circ\varphi(t)$. Since ${\rm image}(\varphi)=l$ never intersects $D$, $\psi(t)\geq 0$ for all $t\in V$. This implies that $\psi$ attains a local minimum at $t=0$. Hence $\psi^{\prime\prime}(0)\geq 0$. Next we want to relate $\psi^{\prime\prime}(0)$ and $\mathcal{H}r(p)$ evaluated along the direction $v$. By an application of the chain rule, we obtain that
$\psi^{\prime}(t)$ $=$ $\frac{\partial r}{\partial z_{1}}(\varphi(t))~v_{1}+\frac{\partial r}{\partial z_{2}}(\varphi(t))~v_{2}+\ldots+\frac{\partial r}{\partial z_{N}}(\varphi(t))~v_{N}$
$\psi^{\prime\prime}(t)$ $=$ $\frac{d}{dt}\left(\frac{\partial r}{\partial z_{1}}(\varphi(t))~v_{1}\right)+\frac{d}{dt}\left(\frac{\partial r}{\partial z_{2}}(\varphi(t))~v_{2}\right)+\ldots+\frac{d}{dt}\left(\frac{\partial r}{\partial z_{N}}(\varphi(t))~v_{N}\right)$
For each $j=1,2,\ldots,N$, we have
$\frac{d}{dt}\left(\frac{\partial r}{\partial z_{j}}(\varphi(t))~v_{j}\right)=\left[\frac{\partial^{2}r}{\partial z_{1}\partial z_{j}}(\varphi(t))~v_{1}+\ldots+\frac{\partial^{2}r}{\partial z_{j}^{2}}(\varphi(t))~v_{j}+\ldots+\frac{\partial^{2}r}{\partial z_{N}\partial z_{j}}(\varphi(t))~v_{N}\right]v_{j}$
Hence, we obtain that $\psi^{\prime\prime}(0)=\displaystyle\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial z_{k}}(p)~v_{j}v_{k}$. The real Hessian $\mathcal{H}r(p)$ evaluated along $v$ is
(5.3) $2\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial\overline{z}_{k}}(p)~v_{j}\overline{v}_{k}+2Re\left(\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial z_{k}}(p)~v_{j}v_{k}\right)$
By the (Levi) pseudoconvexity of $D$ at $p$, we then have for every vector $v\in T_{p}^{\mathbb{C}}(\partial D)$, that:
$\sum_{j,k=1}^{N}\frac{\partial^{2}r}{\partial z_{j}\partial\overline{z}_{k}}(p)~v_{j}\overline{v}_{k}\geq 0.$
The second sum in (5.3) is non-negative, and it equals $2\psi^{\prime\prime}(0)$, which was noted to be non-negative. It therefore follows from (5.3) that the real Hessian $\mathcal{H}r(p)$ along $v$ is non-negative. ∎
Definition 5.9.
Let $D$ be a domain in $\mathbb{C}^{N}$. We say that $D$ admits retracts along an (or, for some, or corresponding to an) open piece of directions at $p\in D$ if there exists a (non-empty) open subset $\Gamma$ of $\partial\mathbb{B}$ such that for each $v\in\Gamma$, there is a retract $Z$ of $D$ passing through $p$ with $v\in T_{0}Z$.
We now restate theorem 1.8 as follows and prove it.
Corollary 5.10.
Consider the balanced pseudoconvex domain given by
$D:=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:|z_{1}|^{q_{1}}+\ldots+|z_{N}|^{q_{N}}<1\}$
where the $q_{j}$’s are all positive numbers less than one. Let $L$ be a linear subspace of $\mathbb{C}^{N}$ which cannot be spanned precisely by vectors only from the standard basis vectors i.e. $L\neq\operatorname{span}(S)$ for any $S\subset B_{std}=\{e_{1},\ldots,e_{N}\}$. Then $D_{L}$ is not a retract of $D$; (thereby, the retracts of $D$ are precisely the orthogonal projections of $D$ onto a span of a finite subset of $B_{std}$). Consequently, $D$ does not admit any retract along any open piece of directions at the origin.
Proof.
Let $C$ denote the convex hull of $\overline{D}$. First we claim that $C=\overline{\mathbb{B}}_{1}$, where $\mathbb{B}_{1}:=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:|z_{1}|+\ldots+|z_{N}|<1\}$ , the $\ell_{1}$-ball in $\mathbb{C}^{N}$. As $q_{j}<1$ for each $j$, we have $|z_{j}|<|z_{j}|^{q_{j}}$. This implies that $\sum_{j=1}^{N}|z_{j}|<\sum_{j=1}^{N}|z_{j}|^{q_{j}}$ and thereby $D\subset\mathbb{B}_{1}$. As $\overline{\mathbb{B}}_{1}$ is convex, $C\subset\overline{\mathbb{B}}_{1}$. Now, $\overline{\mathbb{B}}_{1}$ is the convex hull of the extreme points of $\overline{\mathbb{B}}_{1}$, but these extreme points belong to $\partial D$. Therefore $C=\overline{\mathbb{B}}_{1}$.
Now, we claim that $\partial C\cap\partial D=\{\lambda e_{i}:\lambda\in\mathbb{C},~|\lambda|=1\}$, where $e_{i}$’s are the standard basis vectors of $\mathbb{C}^{N}$. To prove this claim by contradiction, assume that there exists $p$ different from the scalar multiple of $e_{i}$ belonging to $\partial C\cap\partial D$. First, consider the case that every component of $p$ is non-zero. In that case, $p$ is a smooth boundary point of $D$. Since $C$ is convex at $p$, there exists a supporting hyperplane $H_{p}$ to $C$ at $p$. Note that the supporting hyperplane at $p$ coincides with the tangent plane at $p$ because $p$ is a smooth boundary point of $D$. As $D\subset C$ and $p\in\partial C\cap\partial D$, $H_{p}$ serves as a supporting hyperplane to $\partial D$ at $p$. Let $l$ be a (real) affine line passing through $p$ in the direction of $v$, which is contained in $H_{p}$. By the above lemma 5.8, the real Hessian $\mathcal{H}r(p)(v)\geq 0$. Since this holds for all vectors $v\in T_{p}(\partial D)$, $\mathcal{H}r(p)$ is positive semi-definite on $T_{p}(\partial D)$.
Now, we compute the real Hessian of $D$ at $p$. Since $D$ is multicircular, after applying an automorphism on $D$, we may assume $p$ is of the form $(x_{1},\ldots,x_{N})$, where $x_{i}$’s are positive. Upto a non-negative scaling factor of $2$, the real Hessian at the point $(x_{1},\ldots,x_{N})$ along the vector $(v_{1},\ldots,v_{N})$ is given by
$\frac{q_{1}^{2}}{4}x_{1}^{q_{1}-2}|v_{1}|^{2}+\frac{q_{2}^{2}}{4}x_{2}^{q_{2}-2}|v_{2}|^{2}+\ldots+\frac{q_{N}^{2}}{4}x_{N}^{q_{N}-2}|v_{N}|^{2}\\ +Re\left(\frac{q_{1}}{2}(\frac{q_{1}}{2}-1)x^{q_{1}-2}v_{1}^{2}+\frac{q_{2}}{2}(\frac{q_{2}}{2}-1)x_{2}^{q_{2}-2}v_{2}^{2}+\ldots+\frac{q_{N}}{2}(\frac{q_{N}}{2}-1)x_{N}^{q_{N}-2}v_{N}^{2}\right)$
When $v_{1},\ldots,v_{N}$ are positive reals, this real Hessian may be written as
$\frac{q_{1}^{2}-q_{1}}{2}x_{1}^{q_{1}-2}|v_{1}|^{2}+\frac{q_{2}^{2}-q_{2}}{2}x_{2}^{q_{2}-2}|v_{2}|^{2}+\ldots+\frac{q_{N}^{2}-q_{N}}{2}x_{N}^{q_{N}-2}|v_{N}|^{2}$
As $q_{i}<1$, this is negative, contradicting the above conclusion that the real Hessian is positive semi-definite on $T_{p}(\partial D)$. This finishes the proof of the claim.
Next, we consider the case that at least one of the components of $p$ is zero. After a permutation of co-ordinates, we may assume that $p$ is of the form $(p_{1},\ldots,p_{M},0,\ldots,0)$ for some $M<N$. Consider the domain $D^{\prime}:=\{(z_{1},\ldots,z_{M})\in\mathbb{C}^{M}:|z_{1}|^{q_{1}}+\ldots+|z_{M}|^{q_{M}}<1\}$. Note that $p^{\prime}:=(p_{1},\ldots,p_{M})$ is a smooth boundary point of $D^{\prime}$ and $D^{\prime}$ is convex at $p^{\prime}$. By using the same argument applied to $D$, we conclude that this case does not arise. By the above theorem 1.10, if the linear subspace $L$ is not spanned only by vectors $e_{i}$ then $D_{L}$ is not a retract of $D$.
To conclude, let $p=(p_{1},\ldots,p_{N})\in\partial D$ with $p_{i}\neq 0$ for all $i$. Note that $p$ does not belong to a non-trivial linear subspace spanned by standard basis vectors. Hence, there does not exist a retract of $D$ passing through the origin in the direction of $p$. Let $S:=\{(z_{1},\ldots,z_{N})\in\partial D:z_{i}\neq 0~\text{for ~all}~i\}$. Note that $S$ is an open dense subset of $\partial D$ and for each $p\in S$, there does not exist a retract of $D$ passing through the origin in the direction of $p$. Since the complement of $S$ in $\partial D$ forms a nowhere dense subset of $\partial D$, there does not exist any retract of $D$ passing through the origin along an open piece of directions. ∎
Theorem 1.10 can be employed to determine retracts of some non-decoupled domains as well. We demonstrate this in the following.
Corollary 5.11.
Consider the non-decoupled (balanced pseudoconvex) domain given by $\Omega:=\{(z_{1},z_{2})\in\mathbb{C}^{2}:|z_{1}|^{2q_{1}}+|z_{2}|^{2q_{2}}+|z_{1}z_{2}|^{2q_{3}}<1\}$ where the $q_{j}$’s are all positive numbers less than $1/2$. Let $L$ be a linear subspace of $\mathbb{C}^{2}$ not spanned by vectors only from the standard basis vectors. Then $\Omega_{L}$ is not a retract of $\Omega$; (thereby, the retracts of $\Omega$ are precisely the orthogonal projections of $D$ onto a span of the standard basis vectors). Consequently, there are no retracts of $\Omega$ passing through the origin along an open piece of directions.
Proof.
First, we compute the real Hessian of the defining function of $\Omega$ at positive real coordinates $(x,y)$. Consider the function
$\varphi(z_{1},z_{2})=|z_{1}|^{2q_{1}}+|z_{2}|^{2q_{2}}+|z_{1}z_{2}|^{2q_{3}}-1=z_{1}^{q_{1}}\overline{z_{1}}^{q_{1}}+z_{2}^{q_{2}}\overline{z_{2}}^{q_{2}}+(z_{1}z_{2})^{q_{3}}(\overline{z_{1}z_{2}})^{q_{3}}-1$
Note that
$H$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}q_{1}^{2}|z_{1}|^{2(q_{1}-1)}+q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}|z_{1}|^{-2}&q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}(z_{1}\overline{z_{2}})^{-1}\\ q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}(\overline{z_{1}}z_{2})^{-1}&q_{2}^{2}|z_{1}|^{2(q_{2}-1)}+q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}|z_{2}|^{-2}\end{bmatrix}$
In convenience, introduce the notation
$\tilde{H}=\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z_{1}^{2}}&\frac{\partial^{2}\varphi}{\partial z_{1}\partial z_{2}}\\ \frac{\partial^{2}\varphi}{\partial z_{2}\partial z_{1}}&\frac{\partial^{2}\varphi}{\partial z_{2}^{2}}\end{bmatrix}$
where
$\frac{\partial^{2}\varphi}{\partial z_{1}^{2}}=q_{1}(q_{1}-1)|z_{1}|^{2q_{1}}z_{1}^{-2}+q_{3}(q_{3}-1)|z_{1}|^{2q_{3}}z_{1}^{-2}|z_{2}|^{2q_{3}},$
$\frac{\partial^{2}\varphi}{\partial z_{1}\partial z_{2}}=\frac{\partial^{2}\varphi}{\partial z_{2}\partial z_{1}}=q_{3}^{2}|z_{1}z_{2}|^{2q_{3}}(z_{1}z_{2})^{-1},$
$\frac{\partial^{2}\varphi}{\partial z_{2}^{2}}=q_{2}(q_{2}-1)|z_{2}|^{2q_{2}}z_{2}^{-2}+q_{3}(q_{3}-1)|z_{1}z_{2}|^{2q_{3}}z_{2}^{-2}.$
Let $v=[v_{1}~v_{2}]$ and $\overline{v}=[\overline{v_{1}}~\overline{v_{2}}]$. Assume that $v_{1},v_{2}$ are real. Upto a non-negative scaling factor of $2$, the real Hessian at the point $(x,y)$ along the vector $v$ is
$\overline{v}Hv^{T}|_{(x,y)}+Re(v\tilde{H}v^{T})_{(x,y)}=\left[q_{1}(2q_{1}-1)x^{2(q_{1}-1)}+q_{3}(2q_{3}-1)x^{2(q_{3}-1)}y^{2q_{3}}\right]|v_{1}|^{2}+\\ \left[q_{2}(2q_{2}-1)y^{2(q_{2}-1)}+q_{3}(2q_{3}-1)x^{2q_{3}}y^{2(q_{3}-1)}\right]|v_{2}|^{2}+4q_{3}^{2}(xy)^{2q_{3}-1}xy$
The complex tangent space at $(x,y)$ is
$\{(v_{1},v_{2}):\left(q_{1}|z_{1}|^{2(q_{1}-1)}\overline{z}_{1}+q_{3}|z_{1}|^{2(q_{3}-1)}|z_{2}|^{q_{3}}\right)v_{1}=-\left(q_{2}|z_{2}|^{2(q_{2}-1)}\overline{z}_{2}+q_{3}|z_{2}|^{2(q_{3}-1)}|z_{1}|^{q_{3}}\right)v_{2}\}$
This shows that every real vector $v=(v_{1},v_{2})$ in $T_{p}(\partial\Omega)$ for $p\in\partial\Omega$ with positive real co-ordinates whose components $v_{1},v_{2}$ are necessarily opposite in sign. Hence we obtain that the real Hessian at the point $(x,y)$ along $v$ is negative.
Next, we claim that $C=\overline{\mathbb{B}}_{1}$, the $\ell_{1}$-ball in $\mathbb{C}^{2}$. Note that for each $j$, $|z_{j}|<|z_{j}|^{q_{j}}$. This implies that $|z_{1}|+|z_{2}|<|z_{1}|^{q_{1}}+|z_{2}|^{q_{2}}<|z_{1}|^{q_{1}}+|z_{2}|^{q_{2}}+|z_{1}z_{2}|^{q_{3}}$ and thereby $\Omega\subset\mathbb{B}_{1}$. As $\overline{\mathbb{B}}_{1}$ is convex, $C\subset\overline{\mathbb{B}}_{1}$. Note that $\overline{\mathbb{B}}_{1}$ is the convex hull of extreme points of $\overline{\mathbb{B}}_{1}$ which belongs to $\partial\Omega$. Therefore $C=\overline{\mathbb{B}}_{1}$. By using the same arguments in the above corollary, we can prove that $\partial C\cap\partial\Omega=\{\lambda e_{i}:\lambda\in\mathbb{C},~|\lambda|=1\}$, where $e_{i}$’s are the standard basis vectors in $\mathbb{C}^{N}$. After applying the above theorem 1.10, we obtain that if the linear subspace $L$ is not spanned only by vectors $e_{i}$, then $\Omega_{L}$ is not a retract of $\Omega$. Hence, there does not exist retracts of $\Omega$ passing through the origin with an open piece of directions. ∎
@@
We now give a direct and more elementary proof for the retracts of the standard $\ell^{q}$-ball $D_{q}$ (where we no longer assume $N=2$ for that would be an oversimplification unlike the matter in the preceding paragraph), it must first be noted that by proposition 1.3, every holomorphic retract of $\ell_{q}$-ball through the origin is actually linear even when $q<1$. So, if $Y_{q}$ is a holomorphic retract of $D_{q}$ through the origin, then there exists a (linear) projection $P:\mathbb{C}^{N}\longrightarrow\mathbb{C}^{N}$ such that $P(D_{q})=Y_{q}$. Of-course, we are only interested in non-trivial retracts, so $Y_{q}$ is of the form $Y_{q}=D_{q}\cap Y$ for some $r$-dimensional $\mathbb{C}$-linear subspace $Y$ of $\mathbb{C}^{N}$. We only need to show that $Y$ cannot be of the form
@@
$t_{j}=\frac{1}{\left(1+\sum_{\begin{subarray}{c}k=r+1\end{subarray}}^{N}|a_{j,k}|^{q}\right)^{1/q}}.$
@@
Since $P(v_{j})\in Y$, there exists scalars $u_{1,j},\ldots,u_{r,j}$ such that
@@
Note that $P(v_{j})\in\overline{D_{q}}$ for all $j$ . So, we get the following inequality for $1\leq j\leq r$,
@@
By reverse triangle inequality for $q^{\text{th}}$ power where $0<q<1$: $|z-w|^{q}\geq\left||z|^{q}-|w|^{q}\right|$ for all $z,w\in\mathbb{C}$, inequality (5.6) becomes
$\sum_{l=r+1}^{N}|a_{j,l}u_{j,l}|^{q}+\sum_{1\leq k\leq r~~k\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{k,l}\right|^{q}+\sum_{s=r+1}^{N}|a_{j,s}|^{q}-\left|\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\sum_{k=1}^{r}a_{j,l}~a_{k,s}u_{k,l}\right|^{q}\leq 0.$
@@
$\sum_{l=r+1}^{N}|a_{j,l}|^{q}\left(1-\sum_{1\leq k\leq r~k\neq j}|u_{k,l}|^{q}-\sum_{m=r+1}^{N}\left|\sum_{k=1}^{r}a_{k,m}u_{k,l}\right|^{q}\right)+\sum_{1\leq m\leq r~~m\neq j}\left|\sum_{l=r+1}^{N}a_{j,l}u_{m,l}\right|^{q}+\\ \sum_{s=r+1}^{N}|a_{j,s}|^{q}-\sum_{s=r+1}^{N}\sum_{l=r+1}^{N}\left|\sum_{k=1}^{r}a_{j,l}~a_{k,s}u_{k,l}\right|^{q}\leq 0.$
@@
For certain reasons below, we explicitly mention about the special case in dimension $2$. Owing to remarks made in the introduction, it follows as an immediate corollary that $D_{q}=\{(z,w)\in\mathbb{C}^{2}:|z|^{q}+|w|^{q}<1\}$ is not biholomorphic to any product domain. A more important relevance of this example is noted next.
@@
We recall from [Vigue] that fixed point sets of holomorphic self-maps of a any given bounded (not necessarily balanced) convex domain $D$ are precisely the same as its retracts. The example $D_{q}$ in $\mathbb{C}^{2}$ above suffices to show that this coincidence fails as soon as we drop convexity. Indeed, for the non-convex (but pseudoconvex) balanced domain $D_{q}$ in $\mathbb{C}^{2}$ as above, the fixed point set of the holomorphic self-map of $D_{q}$ as simple as the flip $(z,w)\to(w,z)$, namely the linear subspace $\{(z,w)\in\mathbb{C}^{2}\;:\;w=z\}$, is not realizable as a holomorphic retract of $D_{q}$.
The $\ell_{q}$-balls dealt with above are bounded by holomorphically extreme real hypersurfaces. We next discuss examples for the ‘other extreme’ as well as some intermediate cases. We begin with an example where we can show existence of higher dimensional retracts; the purpose being: to record atleast one simple example to affirmatively answer the question as to whether there exists non-homogeneous pseudoconvex domains in $\mathbb{C}^{N}$ (i.e., when $N$ is not restricted to be $2$ or the domain being convex) with the property that there are retracts through every point of the domain and all possible dimensions; given Lempert’s apt remarks about the non-existence of retracts of dimension any bigger than one, in general. As our purpose is in recording the simplest cases first, the next couple of examples may well not be impressive.